OpenStudy (anonymous):

3x-1/3 - x-3/15 = 2x+3/2 (fractions)

hero (hero):

hannah, your parentheses

hero (hero):

\(\large\frac{3x-1}{3} - \frac{x-3}{15} = \frac{2x+3}{2}\)

hero (hero):

Hint: Multiply the first fraction by 5/5

OpenStudy (anonymous):

@Hero why multiply by 5/5?

OpenStudy (anonymous):

i thought tht u would need a common denominator?

OpenStudy (anonymous):

@hannah2013 , you're right, and the best way to do this is by multiplying individual fractions by the LCM of 2, 3 and 15 which is?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

right, now multiply each fraction by 30 and type resulting equation

OpenStudy (anonymous):

(30x-1)/30 (2x-6)/30 (30x+45)/30

hero (hero):

If you multiply the first fraction by 5/5 you get \(\large\frac{15x - 5}{15} - \frac{x - 3}{15} = \frac{2x+3}{2} \)

hero (hero):

Now you can combine the first and second fractions on the left side to get: \(\large\frac{14x - 8}{15} = \frac{2x+3}{2}\) Now cross multiply

OpenStudy (anonymous):

@Hero thts what i got but what about the last fraction?

hero (hero):

2(14x - 8) = 15(2x + 3)

hero (hero):

28x - 16 = 30x + 45

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (anonymous):

 @Hero the idea is to eliminate the denominator completely so that you just work with whole numbers. @hannah2013 , do it like this and write resulting equation